# Both Text and Images Leaked! A Systematic Analysis of Data Contamination in Multimodal LLM

### v2

@@ Abstract
The rapid progression of multimodal large language models (MLLMs) has demonstrated superior performance on various multimodal benchmarks. However, the issue of data contamination during training creates challenges in performance evaluation and comparison. While numerous methods exist for detecting models’ contamination in large language models (LLMs), they are less effective for MLLMs due to their various modalities and multiple training phases. In this study, we introduce a multimodal data contamination detection framework, MM-Detect, designed for MLLMs. Our experimental results indicate that MM-Detect is quite effective and sensitive in identifying varying degrees of contamination, and can highlight significant performance improvements due to the leakage of multimodal benchmark training sets. Furthermore, we explore whether the contamination originates from the base LLMs used by MLLMs or the multimodal training phase, providing new insights into the stages at which contamination may be introduced.
{CJK}
UTF8gbsn
@@ 1 Introduction
The development of MLLMs has exceeded expectations (Liu et al., 2023a; Lin et al., 2023), showcasing extraordinary performance on various multimodal benchmarks (Lu et al., 2022; Liu et al., 2023b; Song et al., 2024), even surpassing human performance. However, due to the partial obscurity associated with MLLMs training (OpenAI, 2023; Reid et al., 2024), it remains challenging to definitively ascertain the impact of training data on model performance, despite some works showing the employment of the training set of certain datasets (Liu et al., 2023a; Chen et al., 2023; Bai et al., 2023b). The issue of data contamination, occurring when training or test data of benchmarks is exposed during the model training phase (Xu et al., 2024), could potentially instigate inequitable performance comparisons among models. This not only creates a dilemma for users in model selection but also poses a significant hurdle to further advancements in this domain.
While numerous works in the field of LLMs have proposed methods for detecting data contamination (Yeom et al., 2018; Deng et al., 2024; Dong et al., 2024), MLLMs, due to their various modalities and multiple training phases (Liu et al., 2023a; Li et al., 2023), face limitations when applying these methods. Therefore, there is a pressing need for a multimodal contamination detection framework specifically tailored for MLLMs.
Figure 1: A description of Multimodal Data Contamination, which can originate from various training stages.
@@
In this study, we carry out a systematic analysis of multimodal data contamination. Initially, we define Multimodal Data Contamination, as it pertains to the modality of data sources exposed to the MLLMs, into two categories: Unimodal Contamination and Cross-modal Contamination, as illustrated in Figure 1. Subsequently, we unveil a detection framework, MM-Detect, which incorporates two methods, Option Order Sensitivity Test and Slot Guessing for Perturbation Caption, designed to handle two common types of Visual Question Answering (VQA) tasks: multiple-choice and caption-based questions, respectively.
@@
To corroborate the validity and sensitivity of our approach, we deliberately induce contamination in MLLMs, thus simulating feasible real-world scenarios. Experimental results indicate that our method proves to be quite effective and sensitive in identifying varying degrees of contamination. Interestingly, our findings reveal that not only does leakage in the multimodal benchmark test set play a role, but the training set can also contribute significantly to enhancing the model’s performance. Then, applying MM-Detect on twelve widely-used MLLMs across five prevalent VQA datasets, we find that both open-source and proprietary MLLMs exhibit contamination with varying degrees.
To further delve into the stage where contamination is introduced, we employ a heuristic method. This method seeks to distinguish whether the contamination originates from the pre-training phase of LLMs or the multimodal training phase. Our findings suggest that the contamination observed in some MLLMs may not necessarily stem from the multimodal training phase. Instead, it could potentially be traced back to the pre-training stage of their respective LLMs.
To the best of our knowledge, our work is the first effort to systematically analyze multimodal data contamination. In conclusion, our research makes several important contributions:
  • We formulate the definition for multimodal contamination and present the MM-Detect framework, comprising two innovative methods specifically designed for effective contamination detection in MLLMs.
  • We demonstrate that leakage from multimodal benchmark data—whether from the training, validation, or test sets—can significantly enhance models’ performance, with this performance gain intensifying as the degree of contamination increases.
  • By employing a heuristic method, we pioneer the exploration into the stage at which contamination is introduced, revealing that it may stem not solely from the multimodal data but could also from the LLMs.
@@ 2 Preliminaries
We formally define the problem of multimodal data contamination and outline the unique challenges associated with its detection.
@@ 2.1 Definition of Multimodal Data Contamination
  • Cross-modal Contamination: The triplet $(x,i,y)$ appears in $D_{\text{vision}}$.
@@ 2.2 Challenges in Multimodal Detection
Figure 2: The overview of proposed MM-Detect framework.
@@
3 MM-Detect
In this section, we introduce the multimodal contamination detection framework, MM-Detect. The core philosophy of MM-Detect is to detect the unusual discrepancies in model performance before and after semantic-irrelevant perturbations. As depicted in Figure 2, this framework operates in two primary steps:
  • The first step is to generate perturbed datasets using two innovative methods: the Option Order Sensitivity Test3.1) and the Slot Guessing for Perturbation Captions3.2), tailored for multiple-choice and image captioning tasks, respectively.
@@ 3.1 Option Order Sensitivity Test
3.2 Slot Guessing for Perturbation Caption
@@
Let $D$ be a dataset containing $n$ datapoints. Each datapoint $d_{i}$ ($i\in\{1,\ldots,n\}$) consists of an image-caption pair, where the caption $S_{i}$ describes the visual features of the corresponding image $I_{i}$. We first apply a back-translation function [^1] We use the Google Translate API for Python to implement back-translation. to $S_{i}$:
@@
$S_{i,\text{mask}}=f_{\text{mask}}(S_{i},K_{i}),~{}sS^{\prime}_{i,\text{mask}}=f_{\text{mask}}(S^{\prime}_{i},K^{\prime}_{i}).$
@@ 3.3 Detection Metrics
Having introduced two detection methods, we now delineate the atomic metrics for the detection pipeline, which consists of two primary steps.
@@
Contamination Degree.
We define the contamination degree solely based on $\Delta$, capturing how much performance changes from $D$ to $D_{\text{pert}}$. The threshold for each severity level (Minor, Partial, Severe) is derived from experimental findings in §4. The entire pipeline, including how these levels are assigned, is summarized in Algorithm 1.
Algorithm 1 MM-Detect Framework
0:  Benchmark dataset $D$, Model $M$
2:  if $D$ is multiple-choice then
3:     Generate perturbed set $D_{\text{pert}}$ via §3.1
4:  else
5:     Generate perturbed set $D_{\text{pert}}$ via §3.2
6:  end if
7:  Compute $CR$, $PCR$, $\Delta$, $\Phi$ using §3.3
8:  if multiple-choice then
10:  else
12:  end if
12:  $CR$, $PCR$, $\Delta$, $\Phi$, $\mathcal{C}$
4 Intentional Contamination
In this section, we address three research questions through intentional contamination:
RQ1: Is MM-Detect an effective detector?
RQ2: How sensitive is MM-Detect?
RQ3: Does training set leakage cause bias?
To answer these questions, we adopt the LLaVA framework and train several 7B-parameter models using intentionally contaminated downstream task data during the visual instruction tuning phase. We then evaluate the degree of contamination in each model. The contamination data are identical to those described in §5.1.
@@ 4.1 MM-Detect is An Effective Detector
Table 1 shows that incorporating contaminated data during training increases both the model’s performance and its sensitivity to perturbations. Compared with the baseline, ScienceQA-contaminated models exhibit average increases in CR and PCR of 9.0% and 5.9%, respectively, while NoCaps-contaminated models show increases of 5.2% and 1.1%. Moreover, all contaminated models demonstrate a marked decrease in $\Delta$, confirming that MM-Detect effectively identifies data contamination. We define the average $\Delta$ across these models as the threshold for severe contamination ($\mathcal{C}_{\text{Severe}}$); specifically, $\Delta\leq-2.9$ indicates severe leakage in the multiple-choice dataset, and $\Delta\leq-5$ indicates severe leakage in the caption dataset.
@@ 4.2 MM-Detect is Sensitive and Fine-grained
As illustrated in Figure 5, increasing contamination from 10% to 50% to 100% results in corresponding increases in CR and PCR, alongside progressively larger $\Delta$ values. We define the $\Delta$ for 50% contamination as the threshold for partial leakage ($\mathcal{C}_{\text{Partial}}$) and that for 10% contamination as the threshold for minor leakage ($\mathcal{C}_{\text{Minor}}$), as detailed in Algorithm 1. These findings confirm that our framework can accurately differentiate between varying leakage levels in benchmark datasets.
4.3 Training Set Leakage Leads to Unfairness
We investigated whether training set leakage biases evaluations by comparing models trained with and without benchmark data contamination. For the ScienceQA experiment, we appended 2000 ScienceQA training samples to the training dataset, creating a contaminated model. For the COCO experiment, we removed the COCO-Caption2017 training data from the original training dataset, resulting in a model without leakage.
@@ 4.3 Training Set Leakage Leads to Unfairness
Table 2 compares the models’ performance. On the ScienceQA test set, the contaminated model outperforms the clean model by 2.9% in CR and 2.3% in PCR, with a $\Delta$ of -0.5, meeting the minor leakage threshold ($\mathcal{C}_{\text{Minor}}$). On the COCO-Caption2017 validation set, the model with COCO data shows a $\Delta$ of -3.2, exceeding the partial leakage threshold ($\mathcal{C}_{\text{Partial}}$). The results indicate that training set leakage inflates performance and that MM-Detect can effectively detect it likewise.
Takeaways Both training and test set leakage can result in unfairness, and the degree of contamination can be detected through MM-Detect effectively.
5 Experiment
@@
Table 3: Comparison of MLLMs on multi-choice datasets. Bold values represent the most significant $\Delta$ or $\Phi$; color codes denote contamination degree: green for $\mathcal{C}_{\text{Minor}}$, yellow for $\mathcal{C}_{\text{Partial}}$, and red for $\mathcal{C}_{\text{Severe}}$.
In this section, we demonstrate the practicality of our methodology in verifying the leakage of multimodal benchmark datasets across several MLLMs.
5.1 Setup
Models.
We conducted extensive evaluations on nine open-source MLLMs, including LLaVA-1.5-7B Liu et al. (2023a), VILA1.5-3B Lin et al. (2023), Qwen-VL-Chat Bai et al. (2023b), fuyu-8b [^3] https://www.adept.ai/blog/fuyu-8b, idefics2-8b Laurençon et al. (2024), Phi-3-vision-128k-instruct Abdin et al. (2024), Yi-VL-6B AI et al. (2024), InternVL2-8B Chen et al. (2023, 2024b), DeepSeek-VL2-Tiny Wu et al. (2024), as well as three proprietary MLLMs: GPT-4o-2024-08-06 OpenAI (2023), Gemini-1.5-Pro-002 Reid et al. (2024), and Claude-3.5-Sonnet-2024-06-20444https://www.anthropic.com/news/claude-3-5-sonnet.
Benchmark Datasets.
5.2 Main Results
@@
Table 4: Comparison of MLLMs on caption datasets.
Multi-choice Datasets.
Table 3 yields several conclusions: (1) Both open-source and proprietary models exhibit contamination. For example, on the ScienceQA training set, both open-source models like LLaVA-1.5-7B and idefics2-8b and proprietary model Gemini-1.5-Pro show minor contamination degree. (2) Proprietary models are more contaminated. Claude-3.5-Sonnet, for instance, registers a severe $\Delta$ with higher $\Phi$ values on both ScienceQA training and test sets, indicating extensive leakage. (3) Training set leakage is more pronounced than test set leakage. On the ScienceQA dataset, models generally exhibit larger $\Delta$ values in the training set, for instance, Claude-3.5-Sonnet shows $\Delta=-5.3$ on training versus $\Delta=-2.4$ on the test set, while most models have near-zero $\Delta$ on the test set. (4) Older benchmarks are more prone to leak. The older ScienceQA test set shows more severe leakage (e.g., Claude-3.5-Sonnet’s $\Delta$ of -2.4) compared to the newer MMStar validation set (e.g., fuyu-8b’s $\Delta$ of -1.2).
@@
6 At Which Stage is Contamination Introduced?
In this section, we investigate the source of contamination in MLLMs. Although the training data for some MLLMs is openly documented, an important question remains: if contamination does not arise during the multimodal training phase, could it stem from the unimodal (pre-training) phase, as defined in §2.1? To address this possibility, we examined the underlying LLMs of the evaluated MLLMs and conducted a series of experiments (§6.1). We also explored the origins of cross-modal contamination arising during visual instruction tuning (§6.2).
6.1 A Heuristic Experiment for Unimodal Contamination Detection
We employed a heuristic approach based on the intuition that if an LLM can correctly answer an image-required question without the image, it may indicate the leakage of that instance.
@@ 6.1 A Heuristic Experiment for Unimodal Contamination Detection
We used MMStar as the benchmark, where every question relies on visual input for correct answers. The tested models include LLaMA2-7B Touvron et al. (2023b) (used by LLaVA-1.5 and VILA), Qwen-7B Bai et al. (2023a) (used by Qwen-VL), Mistral-7B-v0.1 Jiang et al. (2023) (used by idefics2), Phi-3-small-128k-instruct Abdin et al. (2024) (used by Phi-3-vision), Yi-6B AI et al. (2024) (used by Yi-VL), and Internlm2-7B Cai et al. (2024) (used by InternVL2). To discourage random guessing, we appended the prompt “If you do not know the answer, output I don’t know” to the instructions. Accuracy, the frequency with which models correctly answer questions without image input, is reported as the primary metric. Note that we did not evaluate Fuyu-8B and proprietary models since their unimodal LLM and training data remain undisclosed.
@@
Table 5: Contamination rates of LLMs used by MLLMs. $\Phi$M denotes the $\Phi$ of the respective MLLMs.
@@
6.2 Tracing Origins: A Review of MLLM’s Visual Instruction Tuning Data

### v3

@@ Abstract
The rapid advancement of multimodal large language models (MLLMs) has significantly enhanced performance across benchmarks. However, data contamination—unintentional memorization of benchmark data during model training—poses critical challenges for fair evaluation. Existing detection methods for unimodal large language models (LLMs) are inadequate for MLLMs due to multimodal data complexity and multi-phase training. We systematically analyze multimodal data contamination using our analytical framework, MM-Detect, which defines two contamination categories—unimodal and cross-modal—and effectively quantifies contamination severity across multiple-choice and caption-based Visual Question Answering tasks. Evaluations on twelve MLLMs and five benchmarks reveal significant contamination, particularly in proprietary models and older benchmarks. Crucially, contamination sometimes originates during unimodal pre-training rather than solely from multimodal fine-tuning. Our insights refine contamination understanding, guiding evaluation practices and improving multimodal model reliability.
@@ 1 Introduction
The development of MLLMs has exceeded expectations (Liu et al., 2023a; Lin et al., 2023), showcasing extraordinary performance on various multimodal benchmarks (Lu et al., 2022; Liu et al., 2023b; Song et al., 2024), even surpassing human performance. However, due to the partial obscurity associated with MLLMs training (OpenAI, 2023; Reid et al., 2024; Huang et al., 2024), it remains challenging to definitively ascertain the impact of training data on model performance, despite some works showing the employment of the training set of certain datasets (Liu et al., 2023a; Chen et al., 2023; Bai et al., 2023b). The issue of data contamination, occurring when training or test data of benchmarks is exposed during the model training phase (Xu et al., 2024), could potentially instigate inequitable performance comparisons among models. This not only creates a dilemma for users in model selection but also poses a significant hurdle to further advancements in this domain.
Existing contamination detection methods only focus on LLMs (Yeom et al., 2018; Deng et al., 2024; Dong et al., 2024), showing limitations when applied to MLLMs, due to their multimodal data complexity and multi-stage training processes (Liu et al., 2023a; Li et al., 2023; Tie et al., 2025). Thus, systematic analytical frameworks tailored explicitly for multimodal contamination are urgently needed.
Figure 1: An analytical breakdown illustrating different forms and origins of multimodal data contamination across distinct training stages of MLLMs.
In this study, we address three key questions:
  • How can we effectively quantify and detect multimodal data contamination?
  • What is the degree of contamination across different MLLMs and benchmark datasets?
  • When is contamination predominantly introduced—during unimodal pre-training or multimodal fine-tuning?
@@
To comprehensively answer these questions, we first define Multimodal Data Contamination, as it pertains to the modality of data sources exposed to the MLLMs, into two categories: Unimodal Contamination and Cross-modal Contamination, as illustrated in Figure 1. Subsequently, we unveil a detection framework designed explicitly as an analytical tool, MM-Detect, which incorporates two methods, Option Order Sensitivity Test and Slot Guessing for Perturbed Caption, designed to handle two common types of Visual Question Answering (VQA) tasks: multiple-choice and caption-based questions, respectively.
@@
To corroborate the validity and sensitivity of our approach, we deliberately induce contamination in MLLMs, simulating realistic contamination scenarios. Experimental results demonstrate the effectiveness of MM-Detect in identifying varying contamination degrees. Our evaluations on twelve widely-used MLLMs across five prevalent VQA datasets reveal significant contamination among both proprietary and open-source models. Critically, contamination is not only prevalent in multimodal training data but also can originates from unimodal pre-training phases, impacting older benchmarks disproportionately.
In summary, this work provides the first systematic analytical examination of multimodal data contamination, making the following explicit analytical contributions:
  • We analytically characterize multimodal contamination into clearly defined unimodal and cross-modal categories, introducing MM-Detect as an essential analytical tool.
  • We systematically quantify how benchmark leakage inflates performance metrics, providing clear insights into dataset and model susceptibility to contamination.
  • We present novel analytical insights indicating that contamination not solely emerges during the multimodal training stage but could also from unimodal pre-training stage, critically refining current understandings of contamination dynamics.
@@ 2 Preliminaries
We formally define the multimodal data contamination and outline the unique challenges associated with its detection.
@@ 2.1 Definition of Multimodal Data Contamination
  • Cross-modal Contamination: The triplet $(x,i,y)$ or the pair $(x,i)$ appears in $D_{\text{vision}}$.
@@ 2.2 Challenges in Multimodal Detection
Figure 2: The overview of proposed $M$-Detect framework.
@@
3 Detection Framework: MM-Detect
We introduce the multimodal contamination detection framework, MM-Detect, designed explicitly to support our systematic analysis of contamination phenomena. The core philosophy of MM-Detect is to detect the unusual discrepancies in model performance before and after semantic-irrelevant perturbations. As depicted in Figure 2, this framework operates in two primary steps:
  • The first step is to generate perturbed datasets using two methods: Option Order Sensitivity Test3.1) and Slot Guessing for Perturbed Captions3.2), tailored for multiple-choice and image captioning tasks, respectively.
@@ 3.1 Option Order Sensitivity Test
3.2 Slot Guessing for Perturbed Caption
@@
Let $D$ be a dataset containing $n$ datapoints. Each datapoint $d_{i}$ ($i\in\{1,\ldots,n\}$) consists of an image-caption pair, where the caption $S_{i}$ describes the visual features of the corresponding image $I_{i}$. We first apply a back-translation function, where we use the Google Translate API for Python to implement back-translation, to $S_{i}$: [^1] A quantitative analysis of the semantic and lexical similarity between the original and back-translated captions is provided in Appendix D.1.
@@
$S_{i,\text{mask}}=f_{\text{mask}}(S_{i},K_{i}),~S^{\prime}_{i,\text{mask}}=f_{\text{mask}}(S^{\prime}_{i},K^{\prime}_{i}).$
@@ 3.3 Detection Metrics
Detection Metrics serve as the core analytical instruments within MM-Detect. Having introduced two detection methods, we now delineate the atomic metrics for the detection pipeline, which consists of two primary steps.
@@
Compared to methods relying solely on accuracy or perplexity, MM-Detect explicitly highlights performance drop after perturbations, preventing exaggeration or underestimation of contamination. Moreover, it offers advantages of lower computational overhead, higher sensitivity, and effective black-box applicability, thus serving as an essential analytical toolkit in our study.
4 Evaluating MM-Detect with Intentional Contamination
This section tackles our first overarching research question — How can we effectively quantify and detect multimodal data contamination? To operationalise this goal, we break RQ1 into three sub‑questions:
SQ1 (Effectiveness) Is MM-Detect able to detect contamination regardless of where it is injected?
SQ2 (Sensitivity) How finely can MM-Detect measure different leakage levels?
SQ3 (Bias Diagnostic) When training-set data leak, can MM-Detect reveal the evaluation bias?
We answer these sub-questions by adopting the LLaVA framework and training a suite of 7B-parameter models with intentionally contaminated data during the visual-instruction tuning phase. The contamination protocol and data split follow §5.1.
@@ 4.1 MM-Detect is An Effective Detector
Table 1 shows that incorporating contaminated data during training increases both the model’s performance and its sensitivity to perturbations. Compared with the baseline, ScienceQA-contaminated models exhibit average increases in CR and PCR of 9.0% and 5.9%, while NoCaps-contaminated models show increases of 5.2% and 1.1%. Moreover, all contaminated models demonstrate a marked decrease in $\Delta$, confirming that MM-Detect effectively identifies data contamination.
@@ 4.2 MM-Detect is Sensitive and Fine-grained
As illustrated in Figure 5, increasing contamination from 10% to 50% to 100% results in corresponding increases in CR and PCR, alongside progressively larger $\Delta$ values. The findings confirm that our framework can accurately differentiate between varying leakage levels in datasets.
4.3 MM-Detect Diagnoses Evaluation Bias from Training-set Leakage
We investigated whether MM-Detect can detect training set leakage by comparing models trained with and without benchmark data contamination. For the ScienceQA experiment, we appended 2000 ScienceQA training samples to the training dataset, creating a contaminated model. For the COCO experiment, we removed the COCO-Caption2017 training data from the original training dataset, resulting in a model without leakage.
@@ 4.3 MM-Detect Diagnoses Evaluation Bias from Training-set Leakage
Table 2 compares the models’ performance. On the ScienceQA test set, the contaminated model outperforms the clean model by 2.9% in CR and 2.3% in PCR, with a $\Delta$ of -0.5. On the COCO-Caption2017 validation set, the model trained with COCO data shows a $\Delta$ of -3.2. The results indicate that training set leakage inflates performance and that MM-Detect effectively detects it.
MM-Detect also shows strong robustness: $\Delta$ is unaffected by prompt changes (Appendix D.3).
Takeaways Both training and test set leakage can result in unfairness, and the degree of contamination can be detected through $M$-Detect effectively.
5 Assessing the Extent of Contamination in MLLMs
In this section, we systematically quantify the extent of contamination across various MLLMs and benchmarks, addressing our second research question — What is the degree of contamination?
5.1 Setup
Models.
We conducted evaluations on nine open-source MLLMs, including LLaVA-1.5-7B Liu et al. (2023a), VILA1.5-3B Lin et al. (2023), Qwen-VL-Chat Bai et al. (2023b), fuyu-8b [^3] https://www.adept.ai/blog/fuyu-8b, idefics2-8b Laurençon et al. (2024), Phi-3-vision-128k-instruct Abdin et al. (2024), Yi-VL-6B AI et al. (2024), InternVL2-8B Chen et al. (2023, 2024b), DeepSeek-VL2-Tiny Wu et al. (2024), as well as three proprietary MLLMs: GPT-4o-2024-08-06 OpenAI (2023), Gemini-1.5-Pro-002 Reid et al. (2024), and Claude-3.5-Sonnet-2024-06-20444https://www.anthropic.com/news/claude-3-5-sonnet.
Benchmark Datasets.
5.2 Main Results
Multi-choice Datasets.
Table 6 yields several conclusions: (1) Both open-source and proprietary models exhibit contamination. For example, on the ScienceQA training set, both open-source models like LLaVA-1.5-7B and idefics2-8b and proprietary Gemini-1.5-Pro show minor contamination degree. (2) Proprietary models are more contaminated. Claude-3.5-Sonnet, for instance, registers a severe $\Delta$ with higher $\Phi$ values on ScienceQA datasets, indicating extensive leakage. (3) Training set leakage is more pronounced than test set leakage. On the ScienceQA dataset, models generally exhibit larger $\Delta$ values in the training set, for instance, Claude-3.5-Sonnet shows $\Delta=-5.3$ on training versus 2.4-2.4 on the test set, while most models have near-zero $\Delta$ on the test set. (4) Older benchmarks are more prone to leak. The older ScienceQA test set shows more leakage compared to the newer MMStar validation set.
@@ 5.2 Main Results
Table 3: Comparison of MLLMs on multi-choice datasets. Bold values represent the most significant $\Delta$ or $\Phi$; color codes denote contamination degree: green for minor leakage, yellow for partial leakage, and red for severe leakage. [^6] Following §4.1: multi-choice leakage levels by $\Delta$—minor ($1.6<\Delta\leq-0.2$), partial ($2.9<\Delta\leq-1.6$), severe ($\Delta\leq-2.9$); caption leakage levels by $\Delta$—minor ($2.4<\Delta\leq-1.1$), partial ($5.0<\Delta\leq-2.4$), severe ($\Delta\leq-5.0$). See Appendix C.
@@
Table 4: Comparison of MLLMs on caption datasets. Bold values represent the most significant $\Delta$ or $\Phi$; color codes denote contamination degree: green for minor leakage, yellow for partial leakage, and red for severe leakage.
@@
6 Identifying the Origin of Contamination in MLLMs
In this section, we address our third research question — When is contamination predominantly introduced? Although the training data for some MLLMs is openly documented, an important question remains: if contamination does not arise during the multimodal training phase, could it stem from the unimodal (pre-training) phase, as defined in §2.1? To address this possibility, we examined the underlying LLMs of the evaluated MLLMs and conducted a series of experiments (§6.1). We also explored the origins of cross-modal contamination arising during visual instruction tuning (§6.2).
6.1 Heuristic Detection of Unimodal Pre-training Contamination
We employed a heuristic approach based on the intuition that if an LLM can correctly answer an image-required question without the image when random guessing is effectively inhibited, it may indicate the leakage of that instance.
@@ 6.1 Heuristic Detection of Unimodal Pre-training Contamination
We used MMStar as the benchmark, where every question relies on visual input for correct answers. The tested models include LLaMA2-7B Touvron et al. (2023b) (used by LLaVA-1.5 and VILA), Qwen-7B Bai et al. (2023a) (used by Qwen-VL), Mistral-7B-v0.1 Jiang et al. (2023) (used by idefics2), Phi-3-small-128k-instruct Abdin et al. (2024) (used by Phi-3-vision), Yi-6B AI et al. (2024) (used by Yi-VL), and Internlm2-7B Cai et al. (2024) (used by InternVL2). To inhibit random guessing, we appended the prompt “If you do not know the answer, output I don’t know” to the instructions. A sanity check in Appendix D.2 confirms that this uncertainty clause effectively suppresses lucky guesses, validating its inclusion in our main protocol. Accuracy — the frequency with which models correctly answer questions without image input — is reported as the primary metric. Note that we did not evaluate Fuyu-8B and proprietary models since their unimodal LLM components and training data remain undisclosed.
@@
[…]
@@
6.2 Analyzing Cross-modal Contamination in Multimodal Fine-tuning